A proton moves across a magnetic field in a circular path of radius 20 cm. If the flux

density of the field is known to be 0.30 T, find the speed of the proton.



1
Expert's answer
2022-03-11T08:33:23-0500

Answer

the speed of the proton

"v=\frac{qBR}{m}\\=\frac{1.6*10^{-19}*0.30*0.20}{6.67*10^{-27}}\\=1.44*10^{6}m/s"